Calculate the entropy change for 200.0 g of N2 gas that is compressed from a pressure of 0.715

atm to a pressure of 1.16 atm at a constant temperature of 30.0°C. Assume that the gas behaves

ideally.


1
Expert's answer
2021-10-04T23:06:02-0400

Number of moles of nitrogen, n="200\/28=7.14"

universal gas constant ="8.314Jmol^{-1}k^{-1}"

"P_{initial}=0.715, P_{final}=1.16"

Entropy change, "\Delta S = nRln(P_{initial}/P_{final})=7.14\times 8.314ln(\frac{0.715}{1.16})=-28.72JK^{-1}"